OpenStudy (anonymous):

Solve the equation. 2x-3y-5z=-16 x __ __ __ y = __ + __ s + __ t z __ __ __

OpenStudy (amoodarya):

2x-3y-5z=-16 2x=3y+5z-16 x=3/2y+5/2z-8 it was solve for x ca you go on ?

OpenStudy (anonymous):

huh?

zepdrix (zepdrix):

I don't understand what the `s` and `t` are. Are you suppose to parameterize this? :\

OpenStudy (anonymous):

it's vector form for linear algebra ;/ i don't get it either lol.. my book doesn't do a good job explaining this either

zepdrix (zepdrix):

Ohh i see :O Darn, my lin algebra class is going a little slower i guess :c i don't quite know this stuff yet.

OpenStudy (anonymous):

x1=3s+2t x2=4s-3t x3=s x4=t

OpenStudy (anonymous):

@radar can you help me please?

OpenStudy (radar):

Sorry but I never had linear algebra. I am in worse shape than zepdrix lol

OpenStudy (anonymous):

haha okay thank you.